# Vamba: Understanding Hour-Long Videos with Hybrid Mamba-Transformers

### v1

@@ 1 Introduction
Assuming a combined input sequence of $M+N$ tokens, where $M$ is the number of video tokens and $N$ is the number of text tokens, we find that $M$ could be at least 100 times larger than $N$ on many long-video tasks ($M\gg N$). Our model can reduce the training/pre-filling computational complexity from $O(d(M+N)^{2})$ to $O(dMN+d^{2}M)$, where $d$ is the hidden dimension, according to Table 1. In practice, this theoretical improvement may not be fully realized due to the hardware under-optimization for Mamba [24]. Nevertheless, we still observe a more than 50% reduction in GPU memory usage and FLOPs/runtime during training and inference for long video inputs, as shown in Figure 1. Vamba can be efficiently trained using 8XA100 GPUs, whereas other efficient video LMMs such as LongVU [54] and LongLLaVA [64] require 64 and 24 GPUs for training, respectively. By performing a two-stage training, our Vamba achieves a 4.3% improvement over the best efficient video LMMs on the challenging hour-scale video understanding benchmark LVBench [63]. On other long video understanding datasets like Video-MME [21], MLVU [76] and LongVideoBench [66], Vamba also achieves top-notch performance.
@@ 4.2 Main Evaluation Results
Based on our ablation study (Table 2), we adopt Model D as our final design and initialize Vamba from Qwen2-VL-7B. We source $\sim$3M image caption data from CC12M and PixelProse for pretraining and use LLaVA-Video-178K [75], the training set for NExT-QA [67], ActivityNet [70] and PerceptionTest [50] and 2M image-text data from LLaVA-OneVision [33] for instruction-tuning. The full implementation details can be found in Appendix 7.
@@ 5 Related Work
Despite significant advances, current state-of-the-art LMMs remain inefficient at handling long-context multimodal inputs due to the quadratic complexity of causal self-attention operations. To mitigate this issue, methods like Flamingo [3], OpenFlamingo [4], Idefics [31] and LLaMA 3.2 [19] interleave gated cross-attention layers with self-attention layers for vision-text modelling. mPlug-Owl3 [68] develops the hyper-attention layer that integrates self-attention with cross-attention to support multimodal understanding. However, these models generally underperform compared to LLaVA-like transformer-based LMMs. In this study, we identify that one possible reason is the lack of vision token updates for cross-attention-based models. Our Vamba addresses this limitation by incorporating additional Mamba layers, which effectively update vision tokens and enhance multimodal integration.
@@
Another research avenue, exemplified by the Mamba [24, 17] model series, focuses on developing linear-time sequence models to reduce the complexity of autoregressive next-token prediction. Mamba-based [78] and Mamba-Transformer hybrid models [39, 22] have achieved notable success in the realm of pure language modelling. However, these efficient architectures have received little attention in the multimodal understanding domain. LongLLaVA [64] took an initial approach by incorporating the hybrid Jamba [39] model as the LMM decoder, yet its performance in long video understanding remains suboptimal. This shortfall is likely due to the absence of a strong hybrid LLM as the base model and the lack of large-scale training on multimodal instruction-following data.

### v2

@@ 1 Introduction
Assuming a combined input sequence of $M+N$ tokens, where $M$ is the number of video tokens and $N$ is the number of text tokens, we find that $M$ could be at least 100 times larger than $N$ on many long video tasks ($M\gg N$). Our model can reduce the training/pre-filling computational complexity from $O(d(M+N)^{2})$ to $O(dMN+d^{2}M)$, where $d$ is the hidden dimension, according to Table 1. In practice, this theoretical improvement may not be fully realized due to the hardware under-optimization for Mamba [24]. Nevertheless, we still observe a more than 50% reduction in GPU memory usage and FLOPs/runtime during training and inference for long video inputs, as shown in Figure 1. Vamba can be efficiently trained using 8$\times$×A800 GPUs, whereas other efficient video LMMs such as LongVU [55] and LongLLaVA [65] require 64 and 24 GPUs for training, respectively. By performing a two-stage training, our Vamba achieves a 4.3% improvement over the best efficient video LMMs on the challenging hour-scale video understanding benchmark LVBench [64]. On other long video understanding datasets like Video-MME [21], MLVU [76] and LongVideoBench [67], Vamba also achieves top-notch performance.
@@ 4.2 Main Evaluation Results
Based on our ablation study (Table 2), we adopt Model D as our final design and initialize Vamba from Qwen2-VL-7B. We source $\sim$3M image caption data for pretraining and $\sim$7M image and video QA data for instruction-tuning. The full implementation details can be found in Appendix 7.
@@ 5 Related Work
Despite significant advances, current state-of-the-art LMMs remain inefficient at handling long-context multimodal inputs due to the quadratic complexity of causal self-attention operations. To mitigate this issue, methods like Flamingo [3], OpenFlamingo [4], Idefics [33] and LLaMA 3.2 [19] interleave gated cross-attention layers with self-attention layers for vision-text modelling. mPlug-Owl3 [69] develops the hyper-attention layer that integrates self-attention with cross-attention to support multimodal understanding. However, these models generally underperform compared to LLaVA-like transformer-based LMMs. In this study, we identify that one possible reason is the lack of vision token updates for cross-attention-based models. Our Vamba addresses this limitation by incorporating additional Mamba layers to update vision tokens.
@@
Another research avenue, exemplified by the Mamba [24, 17] model series, focuses on developing linear-time sequence models to reduce complexity. Mamba-based [78] and Mamba-Transformer hybrid models [41, 22] have achieved notable success in the realm of pure language modelling. However, these efficient architectures have received little attention in the multimodal understanding domain. LongLLaVA [65] took an initial approach by incorporating the hybrid Jamba [41] model as the LMM decoder, yet its performance in long video understanding remains suboptimal. This shortfall is likely due to the absence of a strong hybrid LLM as the base model and the lack of large-scale training on multimodal instruction-following data.
@@ 8.2 Evaluation Benchmarks
9 Comparison with Contemporary Work
Several contemporary works also investigate hybrid Mamba-Transformer models for long video understanding. For example, STORM [31] and BIMBA [28] utilize Mamba blocks between the vision encoder and LMM decoder as additional processing and compression modules for video tokens, achieving high performance in long video understanding. However, different from Vamba, the overall architecture of the LMM remains unchanged in these methods, with the decoder still relying on full causal self-attention layers for both text and video tokens. As a result, the model architectures proposed in these methods offer limited gains in training and inference efficiency, with any speedup in video processing still primarily attributed to token reduction. In comparison, Vamba directly employs Mamba-2 layers in the LMM decoder and bypasses the self-attention updates for video tokens. This design enables highly efficient video processing even without reducing the number of tokens.
Table 6: Quantitative results for Vamba with token reduction.
Models GPU Mem (MB) LVBench VideoMME MLVU
Vamba 45791 42.4 57.4 65.9
Vamba-TR 33847 41.6 56.9 66.5
10 Combining Vamba and Token Reduction
As mentioned in the paper, we expect Vamba to be compatible with token reduction, and combining Vamba and token reduction can potentially result in similar performance and even higher efficiency. We provide some preliminary results for combining Vamba and token reduction in this section. As shown in Table 6, we can simply uniformly drop 50% of the video tokens during inference (denoted as Vamba-TR) and achieve little performance drop across multiple benchmarks with better efficiency. We believe finetuning Vamba with token reduction can further preserve its capacity and leave this as a future work.